OpenStudy (anonymous):

What is the greatest common factor of 19x7 and 3x5?

OpenStudy (anonymous):

do you mean 19^6 and 3^5

OpenStudy (anonymous):

yes @MIS08

OpenStudy (anonymous):

the answer is x^5

OpenStudy (anonymous):

20x3 + 12x2 + 5x + 3 ? @MIS08

OpenStudy (anonymous):

factor?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Are you giving medals also?

OpenStudy (anonymous):

what are medals ?

OpenStudy (anonymous):

(4x^2+1)(5x+3)

OpenStudy (anonymous):

Medals are the best responses.

OpenStudy (anonymous):

okay ill give you medals 10x3 + 2x2 + 5x + 1? whats this one

OpenStudy (anonymous):

(2x^2+1)(5x+1)

OpenStudy (anonymous):

3x3 + 9x2 + x + 3

OpenStudy (anonymous):

(3x^2+1)(x+3)

OpenStudy (anonymous):

4x3 + 12x2 + 3x + 9

OpenStudy (anonymous):

(4x^2+3)(x+3)

OpenStudy (anonymous):

x3 + 2x2 + x + 2

OpenStudy (anonymous):

(x^2+1)(x+2)

OpenStudy (anonymous):

Find the difference. (6x3 + 7x + 12) - (3x2 + 8x + 8)

OpenStudy (anonymous):

6x^3-3x^2-x+4

OpenStudy (anonymous):

(7x2 + x + 9) - (3x2 + 4x - 5)

OpenStudy (anonymous):

4x^2-3x+14

OpenStudy (anonymous):

(3 - 13x - 7x2) - (5x2 + 12x - 10)

OpenStudy (anonymous):

-12x^2-25x+13

OpenStudy (anonymous):

(9x2 + 10x + 4) - (9x2 + 5x - 1)

OpenStudy (anonymous):

How many more questions do you have?

OpenStudy (anonymous):

ALOT!

OpenStudy (anonymous):

5x+5

OpenStudy (anonymous):

Do you do online school?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Makes sense. So do i.

OpenStudy (anonymous):

how are you so good in math ?

OpenStudy (anonymous):

Lots of studying

OpenStudy (anonymous):

What grade math is this?

OpenStudy (anonymous):

11th.. Solve for x. x2 - 2x = 0

OpenStudy (anonymous):

x=0,2

OpenStudy (anonymous):

x2 + 9x = 0

OpenStudy (anonymous):

Factor? Or what

OpenStudy (anonymous):

solving for x

OpenStudy (anonymous):

x=0,-9

OpenStudy (anonymous):

Solve for x. x2 - 7x = 0

OpenStudy (anonymous):

x=0,7

OpenStudy (anonymous):

Solve for x. x2 + 6x = 0

OpenStudy (anonymous):

x=0,-6

OpenStudy (anonymous):

Solve for x. 2x2 - 4x = 0

OpenStudy (anonymous):

x=0,2

OpenStudy (anonymous):

What are the solutions of the quadratic equation below? -7x2 - 23x + 10 = 0

OpenStudy (anonymous):

do you have multiple choice on this one?

OpenStudy (anonymous):

brb.. inbox me ..